# Benchmarking Multi-Agent Preference-based Reinforcement Learning for Human-AI Teaming

### v1

@@ Abstract
Preference-based Reinforcement Learning (PbRL) is an active area of research, and has made significant strides in single-agent actor and in observer human-in-the-loop scenarios. However, its application within the co-operative multi-agent RL frameworks, where humans actively participate and express preferences for agent behavior, remains largely uncharted. We consider a two-agent (Human-AI) cooperative setup where both the agents are rewarded according to human’s reward function for the team. However, the agent does not have access to it, and instead, utilizes preference-based queries to elicit its objectives and human’s preferences for the robot in the human-robot team. We introduce the notion of Human-Flexibility, i.e. whether the human partner is amenable to multiple team strategies, with a special case being Specified Orchestration where the human has a single team policy in mind (most constrained case). We propose a suite of domains to study PbRL for Human-AI cooperative setup which explicitly require forced cooperation. Adapting state-of-the-art single-agent PbRL algorithms to our two-agent setting, we conduct a comprehensive benchmarking study across our domain suite. Our findings highlight the challenges associated with high degree of Human-Flexibility and the limited access to the human’s envisioned policy in PbRL for Human-AI cooperation. Notably, we observe that PbRL algorithms exhibit effective performance exclusively in the case of Specified Orchestration which can be seen as an upper bound PbRL performance for future research.
Keywords Multi-Agent  $\cdot$ Human-AI Teaming  $\cdot$ Preference-based Reinforcement Learning
@@ 1 Introduction
Modeling co-operation for Human-AI Teaming scenarios has been a long-standing research challenge. With the increase in the pivotal roles AI agents play in day-to-day human activities [1, 2, 3], modeling agents with teaming capabilities from the human-in-the-loop’s perspective has become all the more relevant. This perspective can include what the human-in-the-loop’s expectations are on the AI agent’s policy, what preferences they have on the AI agent’s behavior, or even what their assumptions are on their own policy. Typically, co-operative multi-agent teaming requires all the acting agents to understand each other’s strategies (via explicit communication or implicit deduction from past behavior) and come up with compatible behavior that boosts team performance, and if possible, achieve their personal goals. In this work, we consider the case of a two-agent (Human-AI) cooperative team. While multi-agent cooperation (and even two-agent cooperation) is hard [4], the problem is even more challenging when one of the agents is a human [5, 6, 7]. In particular, it becomes crucial for the AI robot to take into account the cognitive limitations of humans, the preferences humans hold concerning the robot’s goals and strategies, and to align its policies with those of humans to ensure compatibility with policies that humans deem achievable.
@@
Preference-based Reinforcement Learning has been used in several complex domains like robotic manipulation and locomotion to incorporate human preferences on AI agent’s behavior [8]. The initial conceptualization of PbRL focuses on single-agent dynamics with a human in the loop observer available to provide feedback on agent queries [9]. In this work, we provide a first exploration of a more general view of PbRL i.e. in human-AI team setting where the human-in-the-loop co-exists in the same environment as the AI agent and performs actions for achieving the team goal.
The fundamental approach to account for human preferences has been to learn a reward function by querying the human-in-the-loop for their feedback over a pair of agent trajectories. Such preference elicitation becomes more challenging at various levels in a multi-agent scenario. For example, multi-agent (with human-in-the-loop) can span multiple humans or multiple AI agents, which can be cooperative, competitive or act as a swarm system. Moreover, taking human preferences into consideration can be from an external human observing a human-AI team or from the human actively participating in the team, etc. Also, the preferences can themselves be about arbitrary aspects of interaction or behavior between the multiple agents. As a first exploration of Preference-based RL in a multi-agent setting, we constrain our problem to a two-agent cooperative Human-AI team where the human-in-the-loop provides their preferences on the team behavior. Other settings can be explored in future works.
@@
In this work, we specifically look at human-AI teaming as a co-operative game referred to as Human-AI Team PbRL, adopted from [10], with the assumption that the human agent defines the team reward and uses it to compute feasible team strategies, but the AI agent does not have access to it. This reward function is representative of the human’s preferences on the joint behavior of the two agents. We propose a suite of domains, comprising of variants of Highway domain [11] and adapted MuJoCo locomotion domains [12], that requires forced cooperation in a two-agent setting with one agent assumed to be a human agent. As a first exploration of PbRL in Human-AI team setup, we adapt existing state-of-the-art single agent PbRL methods (i.e. RUNE [13], SURF [14] and PEBBLE [15]). Operating under this setting, we argue along two aspects particular of human-AI teaming : Human Flexibility and Access to Human Policy. Human Flexibility, informally, describes how rigid the human agent is with respect to using a specific team strategy. Since the human is acting in the shared world towards a team goal, the robot must find a policy for itself which belongs to set of team strategies considered feasible by the human agent. Otherwise, either the human agent would have to adapt (which can be costlier than having an AI agent adapt) or the team would suffer from sub-optimal performance. Furthermore, the size of this feasible set can be very small or even singleton owing to cognitive limitations of human in computing several strategies or simply due to their preference over team behavior. Another important aspect is the agent’s access to human policy during learning from human preferences. If the agent has zero access to the human’s policy (as selected by the human agent to execute), it has an additional challenge of ‘imagining’ human actions when it queries the human for their preference on the team behavior. On the other hand, the problem can be simplified to a single-agent PbRL setting if the agent has complete knowledge of the human policy and that the human has a single policy in their set of feasible team strategies. We call this as Specified Orchestration, which requires maximal information regarding human behavior and arguably the simplest case for the AI agent, therefore, can be treated as a loose upper bound to the performance of Human-AI PbRL algorithms.
We summarize the contributions of this work as follows:
@@
    This work serves as a first exploration of Preference-based Reinforcement Learning beyond a single-agent setting.
@@
    We propose a suite of two-agent domains which require forced cooperation between the two agents to achieve the team objective. The domains are packaged with human’s feasible set of policies (computed through an oracular approach) and can be used for future research in Human-AI team PbRL.
@@
    We adopt and modify state-of-the-art single-agent PbRL algorithms for our setting and benchmark their performance in policy-learning on the proposed domains.
  4. 4.
    We consider specific aspects of Human-AI team affecting Human-AI PbRL, i.e. Human flexibility in opting for a team strategy and Agent’s access to human Policy, and perform a comprehensive analysis. In particular, we find the case of Specified Orchestration (where the agent has complete access to human’s selected policy and human has a singleton set of team strategies in mind) as a useful upper-bound where the PbRL algorithms exhibit effective performance.
We begin with discussing an example of active human-AI teaming with required co-operation, and analyze the assumptions on the flexibility of the human-in-the-loop in Section 2. Next, we give an overview of the relevant literature in Section 3, followed by describing the preliminaries for this work in Section 4. Section 5 provides a formal framework under which we operate, and defines the notions of Human Flexibility, Agent’s access to Human policy and Specified Orchestration. We describe the proposed domain suite in Section 6.1 and provide an empirical investigation of PbRL for the Human-AI team setup in Section 6. Finally, we conclude with our discussion on the learnings from this work motivating PbRL for human-AI teaming problems in Section 7. An appendix has also been attached as supplementary that includes additional details on the experiments and results, for the ease of reproducibility of our work for future research.
2 Active Human-AI Teaming Example
Figure 1: Two Arm Peg-In-Hole Environment: A representative example for Human-AI Team PbRL Cooperation Game.
Let us consider the example of the Baxter robot, a classic bi-manual robot as shown in Figure 1. There are two robotic arms where one is holding a peg while the other is holding a square plate with a hole in it. Assume that the arm holding the peg is being controlled by the human agent. The human agent has already decided on a policy that they will be moving the peg to a specified location, and insert the peg according to their policy. While there may exist multiple corresponding robot policies that successfully allow for team success, the human agent has specific preferences on the robot behavior as well. For example, the human wants the robot to hold the plate horizontally, reach the location where the peg would be in time and the robot arm should slowly approach that location for safety. The joint task for the two agents is to place the peg in the hole (square plate) successfully without collisions in accordance with the human preference.
The agent is unaware of its goals (holding the plate horizontal at some location), the human preference on how to achieve that goal (manner of reaching the location) and the human partner policy (where the location of peg would be and the trajectory of the human controlled arm to avoid collisions). Therefore, it must utilize some form of communication to gather information about these aspects. Such communication may happen through explicit language-based interaction (which may require an expert understanding of the internal working of the robot, and a robust method for reward function specification that is not prone to reward-hacking), through demonstrations (but in this scenario the human agent holds a peg whereas the robot agent holds the plate which makes providing a demonstration to teach the robot non-trivial) or through binary feedback on agent’s queries (PbRL). In this work we explore the latter case of studying PbRL, where the agent generates behavior trajectories with imagined actions of the human and itself and queries a pair of these behaviors for human feedback. The human agent views the queried trajectory (typically presented as a video) pair and specifies which they prefer more, i.e. which pair achieves the team objective according to them. Recall that the human agent has pre-computed a policy for themselves and has an expectation from the AI partner to complete any remainder objectives, therefore the preference labels are given according to how well the AI-agent is conforming to the human-expected policy. Naturally, the key challenges for the AI agent are to realize the human partner’s chosen policy (such that the queried behaviors show reasonable imagined human agent actions), and how large the team-behavior solution set is based on the human’s internal reward function (for example, the human may find it acceptable to have the plate be placed somewhere in a large region at any angle).
@@ 3 Related Work
Preference-based Reinforcement Learning: Preference-based RL has become a popular paradigm for human-in-the-loop settings where the human is queried for binary feedback by the AI agent to learn a reward function [16, 17, 18, 19, 20, 21, 22]. These methods have not only shown to be effective in teaching an AI agent to accomplish a task where rewards may be otherwise hard to specify such as in the case of tacit knowledge tasks, but also to incorporate the human-in-the-loop’s preferences on how the task needs to be achieved [17]. However, PbRL methods have only been seen in single-agent settings where the AI agent is acting in the environment, and the human is only interacting with the system as and when asked for query feedback [23, 24, 25, 26]. Having realized the potential of PbRL frameworks, we see its applicability in human-AI teaming scenarios where both the human and the AI agent act in the environment to accomplish a team task, and the human may want to specify their preferences through binary feedback distilled into the agent’s reward function. Closest to our research are prior works in multi-agent Inverse Reinforcement Learning (IRL) [27, 28, 29]. We note that IRL and PbRL have key distinctions in assuming expertise of the human in the loop and the modality of the feedback. [30] were one of the first to tackle the multi-agent IRL problem with a central controller. Several works have attempted to study multi-agent IRL for homogeneous agent systems [31, 32, 33, 34] whereas our focus is on a heterogeneous system with a two-player Human-AI cooperative team. More recently, [10] proposed the Cooperative Inverse RL formulation to highlight the limitation of classic formulation which does not allow humans to provide teaching demonstrations. In contrast, our focus is on allowing the AI agent to make explicit queries to the human agent to elicit preferred behaviors. Further, while IRL based setups are dependent on expert-demonstrations by the human agent, such demonstrations may be infeasible (in our Peg-In-Hole example, the human does not control the arm with the square plate), or the human maybe incapable. PbRL, on the other hand, makes it easier for the human agent to provide a binary feedback over trajectories generated by the agent.
Multi-Agent Teaming : Deep Reinforcement Learning has also been used in multi-agent settings [35, 36, 37] such as non-competitive games [38, 39], zero-sum [34], general-sum games [33, 40] and other collaborative environments where rewards are not shared across agents [41, 42]. More recently several works have investigated the challenges in Human-AI teams but commonly assume robot’s access to the team reward function [6, 43, 44, 45]. Prior works have also studied joint-decision making in a two-agent collaboration setting [46, 47, 48, 49] but do not consider inverse RL or PbRL setting.
@@ 4 Preliminaries
Preference-based Reinforcement Learning : In traditional PbRL setups, an agent $A$ interacts with the environment $\mathcal{E}$ with a missing reward function $\mathcal{R}$. It assumes the underlying problem as a MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$. The agent $A$ may not have access to this reward system $\mathcal{R}$ as it may either be hidden [50], hard to specify as seen in tacit-control tasks [17], or otherwise may not be able to capture the goal completely as seen in human-in-the-loop scenarios [51].
The problem of learning the unknown reward function $r_{\psi}$, in single-agent settings, is formulated as a supervised learning problem [17]. A classifier is trained on limited human feedback over agent trajectories. Formally, each agent trajectory is a segment $\sigma$, a sequence of agent observations and actions $\{(s_{k},a_{k}),(s_{k+1},a_{k+1}),\cdot\cdot\cdot(s_{k+N-1},a_{k+N-1})\}$ where $N$ represents the horizon. The human-in-the-loop, i.e., the teacher, gives a feedback $y\in\{0,1\}$ to the agent when queried with a pair of trajectories ($\sigma_{1},\sigma_{2}$), and this feedback is stored as a label along with the two segments in a trajectory buffer $\mathcal{D}$ as ($\sigma_{1},\sigma_{2},y$). Following the Bradley-Terry preference model [52], the preference predictor using the reward function $r_{\psi}$ is learned:
@@
where, $[\sigma_{0}\succ\sigma_{1}]$ implies that trajectory segment $\sigma_{0}$ is preferred over trajectory segment $\sigma_{1}$ by the human teacher. Given the trajectory buffer $\mathcal{D}$, the reward function $r_{\psi}$ is learned by training on minimizing the following Cross-Entropy loss:
$\mathcal{L}_{CE}=-\displaystyle\mathop{\mathbb{E}}_{\begin{subarray}{c}(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}\end{subarray}}\Bigr{[}y(0)\text{log}P_{\psi}[\sigma_{0}\succ\sigma_{1}]+y(1)\text{log}P_{\psi}[\sigma_{1}\succ\sigma_{0}]\Bigr{]}$ (2)
5 Preference-based RL for Human-AI Teaming
Preference based RL has been predominantly studied in single-actor human-observer setting where the human provides feedback on the agent’s behavior but does not act in the shared environment. The aim of this work is to explore PbRL for human-AI team scenario. As discussed in Section 1, there can be multiple interpretations of multi-agent (or a two-agent human-AI team) PbRL. We consider the setup where the human and the AI agents are a cooperative team and the role of the robot is to achieve goals as set by the human agent. Having no knowledge of its objectives, the robot relies on a reward learning process through Preference-based Reinforcement Learning to elicit and learn about the human preferences over the team behavior. We formalize our Human-AI Team PbRL as a special case of the Cooperative Inverse Reinforcement Learning formulation [10] in section 5.
Further, we discuss the two critical aspects for Human-AI team i.e. Human Flexibility over team behavior 5.2 and assumption on agent’s access to human policy 5.3. We discuss a special case with assumed knowledge about human’s policy and the fact that human admits a single team behavior as an optimal solution, referred to as Specified Orchestration. Since we adapt PbRL methods developed for single-actor human-observer setups for our first exploration of Human-AI team PbRL, we discuss the design decisions in Section 6.2.
5.1 Human-AI PbRL Cooperation Game
[10] introduces the problem of Cooperative Inverse Reinforcement Learning but studies it for Apprenticeship Learning. For our Human-AI Team PbRL setting, we modify their formulation as follows:
Definition 5.1.
A Human-AI PbRL cooperation game, as a special case of a Cooperative Inverse Reinforcement Learning game, $M$ is a two-player Markov game with identical payoffs between a human $H$ and a robot $R$. The game is defined as a tuple $M=(S,\{{\mathcal{A}^{H},\mathcal{A}^{R}}\},T(\cdot|\cdot,\cdot,\cdot),R_{\theta}(\cdot,\cdot,\cdot),\gamma)$ with the following definitions:

### v2

# Incorporating Human Flexibility through Reward Preferences in Human-AI Teaming

@@ Abstract
Preference-based Reinforcement Learning (PbRL) has made significant strides in single-agent settings, but has not been studied for multi-agent frameworks. On the other hand, modeling cooperation between multiple agents, specifically, Human-AI Teaming settings while ensuring successful task completion is a challenging problem. To this end, we perform the first investigation of multi-agent PbRL by extending single-agent PbRL to the two-agent teaming settings and formulate it as a Human-AI PbRL Cooperation Game, where the RL agent queries the human-in-the-loop to elicit task objective and human’s preferences on the joint team behavior. Under this game formulation, we first introduce the notion of Human Flexibility to evaluate team performance based on if humans prefer to follow a fixed policy or adapt to the RL agent on the fly. Secondly, we study the RL agent’s varying access to the human policy. We highlight a special case along these two dimensions, which we call Specified Orchestration, where the human is least flexible and agent has complete access to human policy. We motivate the need for taking Human Flexibility into account and the usefulness of Specified Orchestration through a gamified user study. We evaluate state-of-the-art PbRL algorithms for Human-AI cooperative setups through robot locomotion based domains that explicitly require forced cooperation. Our findings highlight the challenges associated with PbRL by varying Human Flexibility and agent’s access to the human policy. Finally, we draw insights from our user study and empirical results, and conclude that Specified Orchestration can be seen as an upper bound PbRL performance for future research in Human-AI teaming scenarios.
@@ 1 Introduction
Figure 1: In this example task, a human and an AI agent need to cook a soup and deliver to the table. Note, that human can either continuously adapt to the AI agent by keeping an expectation on its actions to complete the sub-task or plating the soup in a bowl efficiently, or if possible, complete it independent of the AI agent which is less cognitively demanding but inefficient. We aim to bridge this gap by investigating Human-Flexibility in Human-AI Teaming and showcase how PbRL can be useful.
@@
Preference-based Reinforcement Learning (PbRL) has been studied in several complex domains like robotic manipulation and locomotion to incorporate human preferences on AI agent’s behavior (Wirth et al. 2017). The initial conceptualization of PbRL has primarily focused on single-agent dynamics with a human-in-the-loop observer available to provide feedback on agent queries (Christiano et al. 2017). While several methods have explored improving reward learning and reducing feedback queries, investigating PbRL for multi-agent scenarios remains uncharted.
Modeling cooperation for Human-AI Teaming scenarios has become all the more relevant with the increase in the pivotal roles AI agents play in day-to-day human activities (Wachter, Mittelstadt, and Floridi 2017; Maedche et al. 2019). Typically, cooperative teaming requires all acting agents to understand each other’s strategies (via explicit communication or implicit deduction from past behavior) and execute compatible behavior boosting team performance (see Figure 1). While multi-agent cooperation is hard (Du 2023), the problem is even more challenging when one of the agents is a human (Gal and Grosz 2022; Carroll et al. 2019). It becomes crucial for the AI agent to consider human’s cognitive limitations, their preferences on the robot’s behavior, and align its policies to be compatible with policies that humans deem achievable. However, as we emphasize later, most of the current approaches to modeling Human-AI teaming leave the burden of adapting on the human (Carroll et al. 2019; Strouse et al. 2021; Yu et al. 2023), which often leads to increased cognitive load.
Specifically, in the case of Human-AI teaming, it is imperative to discuss, both, 1) the human-in-the-loop’s expectations or preferences on the joint team behavior, and 2) their constraints or assumptions on their own policy (or policies). To this end, we provide a first exploration of a more general perspective of PbRL extending to the two-agent, i.e., the human-AI team setting where the human-in-the-loop co-exists in the same environment as the AI agent and jointly performs actions for achieving the team goal. We formulate this problem as a Human-AI PbRL Cooperation Game, where the human’s feedback can act as a mechanism for the AI agent to elicit task objectives and human’s preferences on the team behavior.
One might argue that there may exist many solution policy pairs to the Human-AI PbRL Cooperation Game, in which the two agents can cooperate with each other leading to successful task completion. However, for a human-AI team setting (such as in Figure 1), it is commonly assumed that training an AI agent to learn a universal best-response policy leads to optimal team performance, which imperatively leads to the human bearing the burden of adaptation (Carroll et al. 2019; Strouse et al. 2021; Yu et al. 2023) (for example, the green path shown in Figure 1). A key aspect of modeling this adaptation from the human’s perspective is to study the spectrum of flexibility to answer if the human partner is willing to execute a fixed policy or is willing to adapt. Hence, we introduce the notion of Human Flexibility which is defined as a subset of the solution set for the Human-AI PbRL Cooperation Game, consisting of policies that the human is willing to execute for the team task, and we motivate this via a gamified user study conducted over 20 lay users on the task shown in Figure 1, drawing a systematic comparison between the two cases. We find that 50% of the participants complete the cooperative task by not choosing to adapt to the AI agent, while 100% task success is seen when they follow a pre-determined fixed policy. On average, participants took significantly more time to execute their next action while adapting to the other agent. We also noted higher levels of frustration and cognitive load when participants did not follow a fixed policy.
@@
Operating under this setting, we also argue along the aspect of Access to Human Policy, which is costly to obtain, but required for learning the reward function and the corresponding optimal agent policy. If the agent has zero access to the human’s policy (as selected by the human agent to execute), it has an additional challenge of ‘imagining’ human’s actions when it queries the human for their preference on the team behavior. On the other hand, the problem can be simplified to a single-agent PbRL setting if the agent has complete knowledge of the human policy and the human has a single policy in their set of feasible team strategies. We call this case as Specified Orchestration, which requires maximal information regarding human behavior and arguably the simplest case for the AI agent, therefore, can be treated as a loose upper bound to the performance of Human-AI PbRL algorithms. Finally, we re-purpose a suite of domains, comprising of variants of the Highway domain (Leurent 2018) and adapted MuJoCo locomotion domains (Towers et al. 2023), that require forced cooperation in a two-agent setting with one agent assumed to be a human agent in our case. We adapt existing state-of-the-art single agent PbRL methods (i.e. RUNE (Liang et al. 2022), SURF (Park et al. 2022) and PEBBLE (Lee, Smith, and Abbeel 2021)) and extend them to this two-agent setting and draw insights that we believe can be useful for future research in Human-AI teaming scenarios. We summarize the contributions of this work as follows:
@@
    As a first exploration of PbRL beyond a single-agent setting, we investigate the multi-agent Human-AI teaming setting, and formulate this problem as a Human-AI PbRL Cooperation Game.
@@
    We introduce the notion of Human Flexibility and conduct a human subject study showcasing how lay users find following a pre-determined rigid policy (less flexible case) much easier than understanding the agent’s behavior and adapting on the fly (more flexible case). We also discuss the possible cases by varying the agent’s Access to Human Policy.
@@
    We extend state-of-the-art PbRL algorithms for our setting, compare their performance over a suite of two-agent domains, and draw insights on the challenges associated with utilizing PbRL for Human-AI cooperation.
@@ 2 Related Work
Preference-based Reinforcement Learning: PbRL has become a popular paradigm for human-in-the-loop settings where the human is queried for binary feedback by the AI agent to learn a reward function (Fürnkranz et al. 2012; Christiano et al. 2017; Ibarz et al. 2018; Bewley and Lecue 2021; Lee et al. 2021). These methods have been effective in teaching an AI agent to accomplish a task where rewards may be otherwise hard to specify, and also to incorporate the human-in-the-loop’s preferences on how the task needs to be achieved (Christiano et al. 2017). However, PbRL methods have only been seen in single-agent settings where the AI agent is acting in the environment, and the human is only interacting with the system as and when asked for query feedback (Lee, Smith, and Abbeel 2021; Park et al. 2022; Liang et al. 2022). Having realized the potential of PbRL frameworks, we see its applicability in human-AI teaming scenarios where both the human and the AI agent act in the environment to accomplish a team task, and the human can specify their preferences through binary feedback distilled into the agent’s reward function. Closest to our research are prior works in multi-agent Inverse RL (Natarajan et al. 2010; Adams, Cody, and Beling 2022; Pynadath and Tambe 2002). We note that IRL and PbRL have key distinctions in assuming expertise of the human-in-the-loop and feedback modality. Several works have attempted to study multi-agent IRL for homogeneous agent systems (Bogert and Doshi 2017; Lin, Adams, and Beling 2019; Lin, Beling, and Cogill 2017) whereas our focus is on a heterogeneous system with a two-player Human-AI cooperative team. (Hadfield-Menell et al. 2016) proposed the Cooperative IRL formulation to highlight the limitations of classic formulation which does not allow humans to provide teaching demonstrations. In contrast, we focus on allowing the AI agent to make explicit queries to the human agent to elicit preferred behaviors. While IRL-based setups are dependent on expert demonstrations by the human (which may not always be feasible), PbRL, on the other hand, makes it easier to provide a binary feedback over agent trajectories.
Multi-Agent Teaming : Deep Reinforcement Learning has also been used in multi-agent settings (Nguyen, Nguyen, and Nahavandi 2020; Du 2023) such as non-competitive games (Lowe et al. 2017; Gupta, Egorov, and Kochenderfer 2017), zero-sum (Lin, Beling, and Cogill 2017), general-sum games (Lin, Adams, and Beling 2019; Lerer and Peysakhovich 2019) and other collaborative environments where rewards are not shared across agents (Leibo et al. 2017). Several works have investigated the challenges in Human-AI teams but commonly assume robot’s access to the team reward function (Carroll et al. 2019; Strouse et al. 2021; Tylkin, Radanovic, and Parkes 2021). Prior works have also studied joint-decision making in a two-agent collaboration setting (Crandall et al. 2018; Dimitrakakis et al. 2017; Radanovic et al. 2019; Nikolaidis et al. 2017) but do not consider inverse RL or PbRL setting.
@@ 3 Preliminaries
Preference-based Reinforcement Learning : In traditional PbRL setups, an agent $A$ interacts with the environment $\mathcal{E}$ with a missing reward function $\mathcal{R}$. It assumes the underlying problem as a MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$. The agent $A$ may not have access to this reward system $\mathcal{R}$ as it may either be hidden (Ng, Russell et al. 2000), hard to specify as seen in tacit-control tasks (Christiano et al. 2017), or otherwise may not be able to capture the goal completely as seen in human-in-the-loop scenarios (Goodrich, Schultz et al. 2008).
The problem of learning the unknown reward function $r_{\psi}$, in single-agent settings, is formulated as a supervised learning problem (Christiano et al. 2017). A classifier is trained on limited human feedback over agent trajectories. Formally, each agent trajectory is a segment $\sigma$, a sequence of agent observations and actions $\{(s_{k},a_{k}),(s_{k+1},a_{k+1}),\cdot\cdot\cdot(s_{k+N-1},a_{k+N-1})\}$ where $N$ represents the horizon. The human-in-the-loop, i.e., the teacher, gives a feedback $y\in\{0,1\}$ to the agent when queried with a pair of trajectories ($\sigma_{1},\sigma_{2}$), and this feedback is stored as a label along with the two segments in a trajectory buffer $\mathcal{D}$ as ($\sigma_{1},\sigma_{2},y$). Following the Bradley-Terry preference model (Bradley and Terry 1952), the preference predictor using the reward function $r_{\psi}$ is learned:
@@
where, $[\sigma_{0}\succ\sigma_{1}]$ implies that trajectory segment $\sigma_{0}$ is preferred over trajectory segment $\sigma_{1}$ by the human teacher. Given the trajectory buffer $\mathcal{D}$, the reward function $r_{\psi}$ is learned by training on minimizing the following Cross-Entropy loss ($\mathcal{L}_{CE}$):
$\begin{split}-\displaystyle\mathop{\mathbb{E}}_{\begin{subarray}{c}(\sigma_{0},\sigma_{1},\\ y)\sim\mathcal{D}\end{subarray}}\Bigr{[}y(0)\text{log}P_{\psi}[\sigma_{0}\succ\sigma_{1}]+y(1)\text{log}P_{\psi}[\sigma_{1}\succ\sigma_{0}]\Bigr{]}\end{split}$ (2)
4 PbRL for Human-AI Teaming
Preference-based RL has been predominantly studied in single-actor human-observer settings where the human provides feedback on the agent’s behavior but does not act in the shared environment. The aim of this work is to explore PbRL for Human-AI Teaming. We consider a setup where the human and the AI agents are a cooperative team and the role of the robot is to achieve goals as set by the human agent. Having no knowledge of its objectives, the robot relies on a reward learning process through PbRL to elicit and learn about the task objective and human preferences over the team behavior. In this section, we formalize our Human-AI Team PbRL as a special case of the CIRL formulation (Hadfield-Menell et al. 2016). Further, we discuss the two critical aspects of Human-AI Teaming, i.e., Human Flexibility - the assumption on the agent’s access to human policy, and a special case of Specified Orchestration.
Human-AI PbRL Cooperation Game
(Hadfield-Menell et al. 2016) introduces the problem of Cooperative Inverse Reinforcement Learning but studies it for Apprenticeship Learning. For our Human-AI Team PbRL setting, we modify their formulation as follows:
Definition 4.1
A Human-AI PbRL Cooperation Game, $M$ is a two-player Markov game with identical payoffs between a human $H$ and a robot $R$. The game is defined as a tuple $M=(S,\{{\mathcal{A}^{H},\mathcal{A}^{R}}\},T(\cdot|\cdot,\cdot,\cdot),R_{\theta}(\cdot,\cdot,\cdot),\gamma,\mathcal{H}_{\Pi},\mathcal{H}_{\mathcal{B}})$ with the following definitions: